SophiaH4602:

Simplify negative 9 over 6 divided by 3 over negative 2.

supie:

\(\LARGE\frac{-9}{6} \div \frac{3}{-2}\) Does it look like this?

SophiaH4602:

yes!

supie:

So basically \(\LARGE\frac{-9}{6} \div \frac{3}{-2}=\frac{(-3)(2)}{(2)(-3)}=\frac{?}{?}\)

supie:

So -3*2=? 2*-3=?

SophiaH4602:

-6/-6

supie:

Yes, and that can be simplified, so\(\LARGE\frac{-6}{-6}=?\)

SophiaH4602:

1

supie:

Exactly

SophiaH4602:

TYSM!

supie:

np! :)

XioGonz:

The number one is so damn exciting <3 "one" YAYYYYYYYYYYYYY

SophiaH4602:

ok thx supie

supie:

np